OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9

OpenStudy (anonymous):

medal awarded to right answer

OpenStudy (callisto):

3x – 4y = 9 –3x + 2y = 9 consider 3x – 4y =9 -(1) and 9 –3x + 2y = 9 -(2) (1)+(2) 3x – 4y + ( 9 –3x + 2y) = 9+9 9-2y = 18 2y = 9 y = 9/2 = 4.5 put y= 4.5 into (1) 3x -4(4.5) = 9 3x = 27 x=9

OpenStudy (callisto):

wait

OpenStudy (anonymous):

x=-9 because and y=-9 because the 3x cancel out. and then you substitute back into the equation the answer to -9

OpenStudy (anonymous):

4.5 isnt an option

OpenStudy (anonymous):

@order thanks

OpenStudy (callisto):

9-2y = 18 2y = -9 y = -9/2 =- 4.5 put y= 4.5 into (1) 3x -4(-4.5) = 9 3x = -9 x=-3

OpenStudy (anonymous):

No problem, but are you sure you understand?

OpenStudy (anonymous):

yea when she said 4.5 i knew it was wrong

OpenStudy (anonymous):

3x-4y=9 + -3x+2y=9 -------------- -2y=18 y=-9

OpenStudy (callisto):

sorry, but your question misled me when you typed . 3x – 4y = 9 –3x + 2y = 9 it should be considered as 3x – 4y = 9 and 9–3x + 2y = 9 if you mean 3x – 4y = 9 and –3x + 2y = 9 you should at least separate them for a longer distance or add a 'and' in between.

OpenStudy (anonymous):

3x-4(-9)=9 3x+36=9 3x=-27 x=-9

OpenStudy (anonymous):

That's fine. I was confused at first too :p